# RAID-Database: human Responses to Affine Image Distortions

### v1

@@ Abstract
Image quality databases are used to train models for predicting subjective human perception. However, most existing databases focus on distortions commonly found in digital media and not in natural conditions. Affine transformations are particularly relevant to study, as they are among the most commonly encountered by human observers in everyday life.
This Data Descriptor presents a set of human responses to suprathreshold affine image transforms (rotation, translation, scaling) and Gaussian noise as convenient reference to compare with previously existing image quality databases. The responses were measured using well established psychophysics: the Maximum Likelihood Difference Scaling method. The set contains responses to 864 distorted images. The experiments involved 105 observers and more than 20000 comparisons of quadruples of images. The quality of the dataset is ensured because (a) it reproduces the classical Piéron’s law, (b) it reproduces classical absolute detection thresholds, and (c) it is consistent with conventional image quality databases but improves them according to Group-MAD experiments.
@@
Experimental data on subjective differences between images is very useful both in engineering and basic science. On the one hand, the evaluation of subjective image quality is a critical task in image processing. Many perceptual metrics used for different tasks are designed and trained to correlate the subjective quality of distorted images with human perception [1].On the other hand, subjective image distortion has also been used to fit human vision models/metrics. In fact, correlation with subjective distortions has been used to evaluate the human nature of artificial vision systems, and there is a debate about the quality of such databases to be used in these scientific applications. As a consequence, databases containing distorted images paired with human subjective quality scores are essential.
Given the engineering origin of the most common databases, they focus on distortions found in digital environments rather than natural ones [2, 3, 4]. In contrast, the usual type of transformations in nature are due to changes in point of view, motion of objects and changes in illumination. These are all affine transforms of the image: rotations, translations and scaling (affine in the spatial domain), or illumination changes (affine in the amplitude of the signal). Contrast variation of noise which is usual in conventional databases [...], and controlled illumination changes which is not that common, but has also been reported, are extremely useful to fit low-level adaptation models. Similarly, the response to geometric affine transforms could be equally useful to improve mid-level models of shape perception. In any case, studying how perceptual metrics respond to these transformations is important to determine whether their behaviour aligns with human perception [Nuria].
@@ Methods
The database contains a total of 888 images, including 24 reference images obtained from the Kodak Lossless True Color Image Suite. and 864 images distorted with four types of distortions: rotation, translation, scaling, and Gaussian noise. Each distortion has nine incremental levels apart from the reference. As shown in Table 1, these distortion levels were selected to ensure that the range of human visual threshold falls within the first levels for each type of distortion. For rotation, the increment is 2 degrees per level, reaching a maximum rotation of 18 degrees. For translation, each step increases by 0.07 degrees (visual angle) horizontally. Scaling is applied in 1% increments, reaching a maximum of 9% of the image size. Finally, Gaussian noise is added in steps of 0.0009, ranging from 0 to a maximum of 0.008. The maximum level of Gaussian noise was chosen to closely match the maximum noise level in the TID2013 database, allowing the results to be comparable and allowing both data sets to be used together.
@@
Experimental measurement
To obtain the responses for each distorted image, The Maximum Likelihood difference scaling method (MLDS) described by Maloney and Yang[6] was used. This method gets a response curve for each reference image and distortion. This means that only images with the same distortion and reference are compared with each other, resulting in a total of 96 different curves (see Figure 3). In each trial, four images were presented simultaneously: two on the left and two on the right (see figure 2). Observers had to select which pair (left or right) had the largest perceptual difference. 210 quadruples were selected to measure each curve. The 10 stimuli were combined in a way that the two images in one pair $(x_{1},x_{2})$ had a lower distortion level than the two images in the other pair $(x_{3},x_{4})$. Also, for a quadruple $(x_{1},x_{2})$ and $(x_{3},x_{4})$, the difference $||x_{1}-x_{2}|-|x_{3}-x_{4}||$ (call it $\Delta_{difference})$ cannot be larger than 7. This decision was made to avoid combinations where the answer is too obvious in order to make the experiment shorter without removing relevant answers. A total of 105 observers, aged 18 to 60 years, participated in these measurements. Observers visited our laboratory and did the measures on the same monitor under controlled conditions as low external illumination and 80 cm of distance to the screen. Each image was presented on the screen with size 10 cm of diameter, which corresponds to 7.125 vision degrees. Each participant provided informed consent prior to participation and received a gift card as compensation.
@@
For the curve fitting, this method estimates weights that predict the observer’s judgment, except for the first and last parameters, which are fixed at 0 and 1, respectively. The method also estimates the internal noise, which is used to scale the curves. Figure 3 illustrates all the response curves for different reference images and separated in distortions. All curves show an increase in the response as the distortion level increases. However, can be visually perceived that the increase for rotation, translation and scaling follow a more linear trend, whereas Gaussian noise exhibits a more saturating behavior as expected [8].
@@
Piéron’s Law. The correlation of reaction time (RT) and difficulty has been shown by many experimental [15–17] and explained by theoretical studies [18, 19]. $\Delta_{difference}$
Figure 6
@@
Figure 6: Left panel, on the X axis, for each quadruple $(x_{1},x_{2})$, $(x_{3},x_{4})$, the difference $||x_{1}-x_{2}|-|x_{3}-x_{4}||$ at this magnitude $\Delta_{difference}$ For each distortion the Average probability to select the “left”(red) option for all the subjects as a function of the $\Delta_{difference}$ , the difference $||x_{1}-x_{2}|-|x_{3}-x_{4}||$
@@ Acknowledgements
This work was partially funded by the Valencian local government (GVA) under the grant CIGE/2022/066 (grupos emergentes), and by the Ministerio de Ciencia e Innovación grants number PID2020-118071GB-I00 and PDC2021-121522-C21.

### v2

@@ Abstract
Image quality databases are used to train models for predicting subjective human perception. However, most existing databases focus on distortions commonly found in digital media and not in natural conditions. Affine transformations are particularly relevant to study, as they are among the most commonly encountered by human observers in everyday life. This Data Descriptor presents a set of human responses to suprathreshold affine image transforms (rotation, translation, scaling) and Gaussian noise as convenient reference to compare with previously existing image quality databases. The responses were measured using well established psychophysics: the Maximum Likelihood Difference Scaling method. The set contains responses to 864 distorted images. The experiments involved 105 observers and more than 20000 comparisons of quadruples of images. The quality of the dataset is ensured because (a) it reproduces the classical Piéron’s law, (b) observers obtain reasonable psychometric functions depending on the stimulation strength. (c) it reproduces classical absolute detection thresholds, and (d) it is consistent with conventional image quality databases but improves them according to Group-MAD experiments.
@@
Experimental data on subjective differences between images is very useful both in engineering and basic science. On the one hand, the evaluation of subjective image quality is a critical task in image processing. Many perceptual metrics used for different tasks are designed and trained to correlate the subjective quality of distorted images with human perception [1].On the other hand, subjective image distortion has also been used to fit human vision models/metrics  [2, 3]. In fact, correlation with subjective distortions has been used to evaluate the human nature of artificial vision systems, and there is a debate about the quality of such databases to be used in these scientific applications  [4]. As a consequence, databases containing distorted images paired with human subjective quality scores are essential.
Given the engineering origin of the most common databases, they focus on distortions found in digital environments rather than natural ones [5, 6, 7, 8]. In contrast, the usual type of transformations in nature are due to changes in point of view, motion of objects and changes in illumination. These are all affine transforms of the image: rotations, translations and scaling (affine in the spatial domain), or illumination changes (affine in the amplitude of the signal). Contrast variation of noise which is usual in conventional databases [...], and controlled illumination changes which is not that common, but has also been reported, are extremely useful to fit low-level adaptation models. Similarly, the response to geometric affine transforms could be equally useful to improve mid-level models of shape perception. In any case, studying how perceptual metrics respond to these transformations is important to determine whether their behaviour aligns with human perception [9].
@@ Methods
The database contains a total of 888 images, including 24 reference images obtained from the Kodak Lossless True Color Image Suite. and 864 images distorted with four types of distortions: rotation, translation, scaling, and Gaussian noise. Each distortion has nine incremental levels apart from the reference. As shown in Table 1, these distortion levels were selected to ensure that the range of human visual threshold falls within the first levels for each type of distortion. For rotation, the increment is 2 degrees per level, reaching a maximum rotation of 18 degrees. For translation, each step increases by 0.07 degrees (visual angle) horizontally. Scaling is applied in 1% increments, reaching a maximum of 9% of the image size. Finally, Gaussian noise is added in steps of 0.0009, ranging from 0 to a maximum of 0.008 in variance. The maximum level of Gaussian noise was chosen to closely match the maximum noise level in the TID2013 database, allowing the results to be comparable and allowing both data sets to be used together.
@@
Experimental measurement of the responses
To obtain the responses for each distorted image, The Maximum Likelihood difference scaling method (MLDS) described by Maloney and Yang[11] was used. This method gets a response curve for each reference image and distortion. This means that only images with the same distortion and reference are compared with each other, resulting in a total of 96 different curves (see Figure 3). In each trial, four images were presented simultaneously: two on the left and two on the right (see figure 2). Observers had to select which pair (left or right) had the largest perceptual difference. 210 quadruples were selected to measure each curve. The 10 stimuli were combined in a way that the two images in one pair $(x_{1},x_{2})$ had a lower distortion level than the two images in the other pair $(x_{3},x_{4})$. Also, for a quadruple $(x_{1},x_{2})$ and $(x_{3},x_{4})$, the difference $||x_{1}-x_{2}|-|x_{3}-x_{4}||$ (call it $\Delta_{\textrm{\emph{difference}}})$ cannot be larger than 7. This decision was made to avoid combinations where the answer is too obvious in order to make the experiment shorter without removing relevant answers. A total of 105 observers, aged 18 to 60 years, participated in these measurements. Observers visited our laboratory and did the measures on the same monitor under controlled conditions as low external illumination and 80 cm of distance to the screen. Each image was presented on the screen with size 10 cm of diameter, which corresponds to 7.125 vision degrees. Each participant provided informed consent prior to participation and received a gift card as compensation.
@@
For the curve fitting, this method estimates weights that predict the observer’s judgment, except for the first and last parameters, which are fixed at 0 and 1, respectively. The method also estimates the internal noise, which is used to scale the curves.
While conducting the MLDS experiments the reaction time was recorded because (1) it is useful to assess the quality of the responses if they are consistent with the classical Pieron’s law, and (2) this behavior may be related to the sensitivity of the observers (or slope of the response curves).
Responses
Figure 3 illustrates all the response curves for different reference images and separated in distortions. All curves show an increase in the response as the distortion level increases. However, can be visually perceived that the increase for rotation, translation and scaling follow a more linear trend, whereas Gaussian noise exhibits a more saturating behavior as expected [12].
@@
The correlation of reaction time (RT) and task difficulty has been shown by many experimental, [15–17] and explained by theoretical studies [18, 19]. The form of Piéron’s Law, applied for each transformations, reflects a power-law relationship between stimulus degree of distortion and reaction time (RT) Fig.6 left. Fig.6 emphasizes that as the distortion becomes stronger, the brain processes it faster, resulting in shorter reaction times. This is thought to arise from the brain’s neural efficiency in encoding higher-intensity signals. As shown in Fig.6-Right, RT decreases as a function of the ease of the trial bigger $\Delta_{diff}$. This V inverted-shaped pattern has been previously associated to decision difficulty [5, 7, 20–22]. Therefore, and since we do not have a direct measure of difficulty for the different transformations,
@@
Figure 6: Left panel, the average probability to select the “left” option for all the subjects as a function of the $\Delta_{diff}$ (the difference $||x_{1}-x_{2}|-|x_{3}-x_{4}||$ for the quadruplets). Right panel: Average Reaction time in seconds (RT) for for all the subjects as a function of the $\Delta_{diff}$ for all the distortion .
@@ Acknowledgements
This work was partially funded by the Valencian local government (GVA) under the grant CIGE/2022/066 (grupos emergentes), and by the Ministerio de Ciencia e Innovación grants number PID2020-118071GB-I00 and PDC2021-121522-C21, and by the grant BBVA Foundations of Science program: Mathematics, Statistics, Computational Sciences and Artificial Intelligence (VIS4NN).